# Generalised Causal Dantzig

### v1

@@ 4 Identifying causal drivers in two empirical studies
Here we illustrate the use of the generalised causal Dantzig method on two empirical studies. In section 4.1 we consider data from a women’s fertility study in the US. The aim is to identify the causal drivers of women’s reproductive fertility. In section 4.2 we consider a socio-economic study for identifying potential causal drivers of high income.
@@ 4.1 Causal determinants of fertility
The National Opinion Resource Center’s General Social Survey, which is currently available for 34 editions between 1972 and 2022 (GSS, 2022), can be used for the identification of the causal determinants of women’s fertility. The survey is conducted on English-speaking women living in non-institutional arrangements in the United States. As in Sander (1992), we consider women between the age of $35$ and $54$, as younger women may not have completed their schooling or their reproductive period. Across the $34$ editions of the survey, this results in a sample size of $n=16,269$ women. The target variable of interest is the “number of children ever born,” whereas the potential causal determinants are investigated among $p=8$ available covariates:
@@
We fit an additive Poisson regression model with all variables, where we consider thin-plate regression splines for the 5 numerical variables (years of schooling, mother’s and father’s years of schooling, year of study, age) and factorial main effects for the remaining 3 categorical variables. The variables produce a relatively good fit, with 8.04% of deviance explained. The results correspond to other association studies conducted in the literature, with years of schooling, age, race, living environment and years of study picked up as the most significant variables (Sander, 1992; Wooldridge, 2009).
@@
Figure 4: Causal determinants of women’s fertility detected by the proposed generalised causal Dantzig framework show how (a) time is associated to a decline in the number of children born; (b) age has a non-linear increasing causal effect on the number of children born; (c) high levels of education cause a sharp drop in fertility; (d) mother’s levels of education also cause a decrease in the number of children born.
In particular, Figure 4(a) shows an overall decline over time in the number of children born, but with an interesting rise between 2000 and 2010. In the context of causality, this variable is most likely capturing unmeasured time-varying causal determinants of fertility. Figure 4(b) shows a non-linear and increasing causal effect of women’s age on the number of children born, which is to be expected. The non-linear association between this variable and the response is picked up also by Wooldridge (2009) who includes a quadratic polynomial term for age in the linear model. It is interesting to note how a search of the causal model among the linear models does not pick up any causal determinant, even when one includes a quadratic term in age. This shows the flexibility of the proposed additive modelling framework in allowing to include smooth effects. Finally, Figure 4(c) shows a marked non-linear causal effect of years of schooling on fertility, with high levels of education causing a sharp drop in fertility. This is pointed out also by Sander (1992). Although less pronounced, the same negative effect is found in relation to the mother’s years of schooling. Among the categorical variables, we find a positive causal effect associated to being black, compared to other races (estimated causal coefficient 0.15) and to living in north-central or west American regions, compared to south (estimated causal coefficients both equal to 0.07). This is also in line with previous studies (Sander, 1992).
@@ 4.2 Causal determinants of income
We perform generalised causal Dantzig among the family of additive logistic regression models, where we consider thin-plate regression splines for the 3 numerical variables (education level, age and hours-per-week) and factorial main effects for the remaining 5 categorical variables. A full search of all 256 models with $\alpha=5\%$ returns a model with age, education level, marital status and occupation as the causal determinants.
Figure 5: Causal determinants of high income detected by the proposed generalised causal Dantzig framework show how (a) age has an increasing causal effect on income, during the early working years; (b) education level has an increasing, close to linear, association with high earning.
Figure 5, for the two numerical variables, shows how age has a non-linear causal effect on income, with a pronounced increasing effect during the early working years, while education level has an increasing, close to linear, association with high earning. As for the categorical variables, being male makes it 73% more likely to be a high earner, whereas being married makes is roughly 7 times more likely than any other marital status level. With respect to type of occupation, white collar jobs, sales jobs and other professionals are between 60% and 160% more likely to be high earners, compared to blue collar jobs, services and other job types.
@@ 5 Conclusion
In this paper, we have proposed a novel semi-parametric approach to causal discovery, when, conditional on its causal parents, the target variable of interest has a distribution in the exponential dispersion family. We have characterised the causal model by means of two properties: invariance of the expected Pearson risk and maximization of the population likelihood. We have proposed a generalised causal Dantzig procedure, that consists of (i) a statistical testing procedure of the empirical Pearson risk, returning a list of candidate causal models; and (ii) a BIC selection step to refine the list of candidate models, as models containing variables that are d-separated from the target by the causal parents lead to the same perfectly dispersed Pearson risk as the causal model.
For generalised linear models with a known dispersion parameter, such as for Poisson or Binomial, the Pearson risk is known. For these settings, the causal model can be recovered with data from a single environment. For other distributional families, the method requires an a priori value for the dispersion parameter. This may be possible in certain applications where the negative Binomial is used. If not available, data from multiple environments can be considered in order to fix the dispersion parameter across these environments. This is the idea of the causal Dantzig estimator in the Gaussian or additive noise setting. Given that our method extends this idea to any exponential family distribution, we refer to it as the generalised causal Dantzig.
@@ Acknowledgments
EW acknowledges support from the Swiss National Science Foundation (grant 192549). VV received financial support from the MUR-PRIN grant P2022BNNEY (CUP E53D23016580001).

### v2

@@ 4 Identifying causal drivers in empirical studies
Here we illustrate the use of the causal GLM method on three empirical studies. In section 4.1, we consider data generated under a controlled experiment. In this case, there is knowledge about the true underlying physical model which can be used for validation. In section 4.2 we consider data from a women’s fertility study in the US. The aim is to identify the causal drivers of women’s reproductive fertility. In section 4.3 we consider a socio-economic study for identifying potential causal drivers of high income.
4.1 Causal chambers: a controlled causal experiment
Gamella et al. (2025) describes a unique experimental program, called causal chambers, in which physical laws are combined with a replicable experimental set-up to generate on-demand datasets that can be used for causal discovery validation. Here, we consider a dataset from a light tunnel set-up. The light tunnel contains a controllable light source and a variety of sensors, which allow to measure the light intensity at different points of the tunnel. Using the notation of Gamella et al. (2025), the following variables are of particular importance:
  • $R$, $G$, $B$: Color setting of the light source
  • $\tilde{C}$: Current drawn by the light source
  • $\tilde{I}_{1}$,$\tilde{I}_{2}$, $\tilde{I}_{3}$: Infrared-intensity measurement at the first, second and third sensor
  • $\tilde{V}_{1}$, $\tilde{V}_{2}$, $\tilde{V}_{3}$: Visible-intensity measurement at the first, second and third sensor
  • $L_{i1}$, $L_{i2}$: Brightness of the two LEDs placed at the $i$th sensor
  • $\theta_{1}$, $\theta_{2}$: Position of the first and second polarizer.
[…]
Figure 7: (a) Causal graph underlying the light tunnel causal chamber experiment, taken from Gamella et al. (2025). (b) Correlation heatmap among the main variables in the system.
We focus on the variable $\tilde{I}_{2}$ in the uniform reference setting. A ground-truth causal graph of the system related to this variable is reported in Figure 7(a) (Gamella et al., 2025). Since in this environment the photodiodes used by the light sensors are set to the large infrared photodiodes ($D^{I}_{2}=2$) and their exposure time is set to the largest value ($T^{I}_{2}=3$), i.e., both are set to a constant, the causal parents of the target variable $\tilde{I}_{2}$ in the causal graph are the light source colors ($R,G,B$) and the brightness of the two LDEs placed at this second sensor ($L_{21}$, $L_{22}$).
As the variable $\tilde{I}_{2}$ is continuous, we construct a binary target variable that indicates if the second light sensor $\tilde{I}_{2}$ is above a certain threshold, which we fix to $2000$. This means that $\tilde{I}_{2}^{\mbox{\scriptsize bin}}$ is Bernoulli distributed with a probability that depends on the distribution $F_{\tilde{I}_{2}}$ of $\tilde{I}_{2}$, which in turn depends on the causal parents, i.e.,
$\left\{\begin{array}[]{l}\tilde{I}_{2}^{\mbox{\scriptsize bin}}\sim\mbox{Bern}(\pi)\\ \pi=F^{-1}_{\tilde{I}_{2}}(2000;R,G,B,L_{21},L_{22})\end{array}\right.$ (2)
The aim is to recover the causal parents of this target variable in the causal graph. As the response is binary, it is a natural choice to use the logistic additive regression version of the proposed causal approach, i.e., for $k=1,\ldots,10,000$,
$\left\{\begin{array}[]{l}\tilde{I}_{2k}^{\mbox{\scriptsize bin}}\sim\mbox{Bern}(\pi_{k})\\ \mbox{logit}~{}\pi_{k}=f_{1}(X_{k1})+\ldots+f_{p}(X_{kp})\end{array}\right.$ (3)
where $X_{1},\ldots,X_{p}$ represent the available variables in the uniform reference dataset provided by Gamella et al. (2025) as the observational environment for the analysis. However, it is important to note that the non-linear additive form in (3) may not be rich enough to capture the true functional form in (2). This may impede the discovery of the true causal parents. Moreover, as can be seen in Figure 7(b), $L_{21}$ and $L_{22}$ appear to be not predictive of $\tilde{I}_{2}^{\mbox{\scriptsize bin}}$ and are not included in the Markov blanket under any analysis for the available number of observations $n$. For this reason, we exclude them from the causal discovery algorithm.
For the causal discovery, we restrict the search to the $p=9$ most predictive variables, found via a generalized additive model with all variables included and a null space penalization (Wood, 2017). In particular, we consider the following potential causal parents: $R$, $G$, $B$, $\tilde{C}$, $\tilde{I}_{1}$, $\tilde{I}_{3}$, $\tilde{V}_{1}$, $\tilde{V}_{2}$, $\tilde{V}_{3}$. An exhaustive search across all $512$ models, using bootstrap for calculation of p-values and setting $\alpha=10\%$ as significance level, returns the causal model with $R$, $G$ and the three visible-intensity measurements as causal parents of $\tilde{I}_{2}^{\mbox{\scriptsize bin}}$. This is only partly reflecting the true generative process, since $B$ has not been detected and the non-causal variables $\tilde{V}_{1}$, $\tilde{V}_{2}$, and $\tilde{V}_{3}$ are most likely compensating (i) for the absence of this causal parent (indeed, they are the three most correlated variables with $B$ in Figure 7(b)) and (ii) for the miss-specification of the true link function in (2) by the additive logistic model in (3).
@@ 4.2 Causal determinants of fertility
The National Opinion Resource Center’s General Social Survey, which is currently available for 34 editions between 1972 and 2022 (GSS, 2022), can be used for the identification of the causal determinants of women’s fertility. The survey is conducted on English-speaking women living in non-institutional arrangements in the United States. As in Sander (1992), we consider women between the age of $35$ and $54$, as younger women may not have completed their schooling or their reproductive period. Across the $34$ editions of the survey, this results in a sample size of $n=16,269$ women. The target variable of interest is the number of children ever born, whereas the potential causal determinants are investigated among $p=8$ available covariates:
@@
We fit an additive Poisson regression model with all variables, where we consider thin-plate regression splines for the 5 numerical variables (years of schooling, mother’s and father’s years of schooling, year of study, age) and factorial main effects for the remaining 3 categorical variables. The variables produce a relatively good fit, with 8.0% of deviance explained. The results correspond to other association studies conducted in the literature, with years of schooling, age, race, living environment and years of study picked up as the most significant variables (Sander, 1992; Wooldridge, 2009).
@@
Figure 8: Causal determinants of women’s fertility detected by the proposed invariant risk framework show how (a) time is associated to a decline in the number of children born; (b) age has a nonlinear increasing causal effect on the number of children born; (c) high levels of education cause a sharp drop in fertility; (d) mother’s levels of education also cause a decrease in the number of children born.
In particular, Figure 8(a) shows an overall decline over time in the number of children born, but with an interesting rise between 2000 and 2010. In the context of causality, this variable is most likely capturing unmeasured time-varying causal determinants of fertility. Figure 8(b) shows a nonlinear and increasing causal effect of women’s age on the number of children born, which is to be expected. The nonlinear association between this variable and the response is picked up also by Wooldridge (2009) who includes a quadratic polynomial term for age in the linear model. It is interesting to note how a search of the causal model among the linear models does not pick up any causal determinant, even when one includes a quadratic term in age. This shows the flexibility of the proposed additive modelling framework in allowing to include smooth effects. Finally, Figure 8(c) shows a marked nonlinear causal effect of years of schooling on fertility, with high levels of education causing a sharp drop in fertility. This is pointed out also by Sander (1992). Although less pronounced, the same negative effect is found in relation to the mother’s years of schooling. Among the categorical variables, we find a positive causal effect associated to being black, compared to other races (estimated causal coefficient 0.15) and to living in north-central or west American regions, compared to south (estimated causal coefficients both equal to 0.07). This is also in line with previous studies (Sander, 1992).
@@ 4.3 Causal determinants of income
We apply the invariant risk method among the family of additive logistic regression models, where we consider thin-plate regression splines for the 3 numerical variables (education level, age and hours-per-week) and factorial main effects for the remaining 5 categorical variables. A full search of all 256 models with $\alpha=5\%$ returns a model with age, education level, marital status and occupation as the causal determinants.
Figure 9: Causal determinants of high income detected by the proposed invariant risk framework show how (a) age has an increasing causal effect on income, during the early working years; (b) education level has an increasing, close to linear, association with high earning.
Figure 9, for the two numerical variables, shows how age has a nonlinear causal effect on income, with a pronounced increasing effect during the early working years, while education level has an increasing, close to linear, association with high earning. As for the categorical variables, being male makes it 73% more likely to be a high earner, whereas being married makes is roughly 7 times more likely than any other marital status level. With respect to type of occupation, white collar jobs, sales jobs and other professionals are between 60% and 160% more likely to be high earners, compared to blue collar jobs, services and other job types.
@@ 5 Conclusion
In this paper, we have proposed a novel semi-parametric approach to causal discovery, when, conditional on its causal parents, the target variable of interest has a distribution in the exponential dispersion family. We have characterized the causal model by means of two properties: invariance of the expected Pearson risk and maximization of the expected likelihood. We have proposed a causal generalized linear model procedure, that consists of (i) a statistical testing procedure of the empirical Pearson risk, returning a list of candidate causal models; and (ii) a BIC selection step to refine the list of candidate models, as models containing variables that are d-separated from the target by the causal parents lead to the same perfectly dispersed Pearson risk as the causal model.
For generalized linear models with a known dispersion parameter, such as for Poisson or Binomial, the Pearson risk is known. For these settings, the causal model can be recovered with data from a single environment. This is innovative in the context of invariant causal prediction, as the existing methods based on this idea require availability of data from multiple distinct experiments. Given the widespread use of Poisson and logistic regression models in association studies, this paper provides the means to tackle these studies from a causal perspective. For other distributional families, the method requires an a-priori value for the dispersion parameter, which may sometimes be available, such as for the negative binomial. If not available, data from multiple environments can be considered in order to fix the dispersion parameter across these environments, in line with similar invariance prediction methods available in the literature also for general response types (Kook et al., 2024).
A further challenge of invariance prediction methods is the computational complexity of searching through the large space of potential causal models. In this paper, we propose a stepwise alternative to the exhaustive search across all models. The simulation studies show a natural reduction in computational cost without compromising accuracy too much. In order to increase computational efficiency, future work is looking at approximations of the test statistic under the null hypothesis, which are currently available only for some generalized linear models. In particular, we have used the fact that the Pearson statistics of a Poisson regression model can be well approximated by a chi-squared distribution under the null hypothesis to avoid time-consuming bootstrap procedures in this case.
@@ Acknowledgments
EW acknowledges support from the Swiss National Science Foundation (grant 192549). VV received financial support from the European Union (Next Generation EU Project PRIN 2022 PNRR, Mission 4 Component 2 Investment 1.1, Protocol Number P2022BNNEY, CUP E53D23016580001).